OpenStudy (anonymous):

In ABC, AB = x, BC = y, and CA = 2x. A similarity transformation with a scale factor of 0.5 maps ABC to MNO, such that vertices M, N, and O correspond to A, B, and C, respectively. If OM = 5, what is AB?

OpenStudy (anonymous):

AB=5, because: MO = 0.5*AC MO = 0.5*2x 5 = 0.5*2x 5 = x x = 5